Prove that for all integer n>=3, P(n+1,3) - P(n,3) = 3P(n,2)

.



1
Expert's answer
2021-10-19T02:04:14-0400

"\\\text {To Prove}: P(n+1 , 3)-P(n, 3)=3 P(n, 2) \quad"

"\\ \text{Proof: LHS}=
\\ [\text{Using}\ P(n, \mu)=\frac{n !}{(n-\mu) !}]
\\ \frac{(n+1) !}{(n+1-3) !}-\frac{n !}{(n-3) !}
\\ = \frac{n !(n+1)}{(n-2) !}-\frac{n !}{(n-3) !}
\\ = \frac{(n+1) n !}{(n-2)(n-3) !}-\frac{n !}{(n-3) !}
\\ = \frac{n !}{(n-3) !}\left[\frac{(n+1)}{(n-2)}-1\right]
\\ = \frac{n !}{(n-3) !}\left[\frac{n+1-n+2}{n-2}\right]
\\ =\frac{n !}{(n-3) !}\left[\frac{3}{(n-2)}\right]
\\ =\frac{3 n !}{(n-2)(n-3) !}=3 \frac{n !}{(n-2) !}
\\ = \quad 3 P(n, 2)= \text{RHS}
\\ \text{Hence proved.}"